# SpecExtend: A Drop-in Enhancement for Speculative Decoding of Long Sequences

### v1

@@ Abstract
Speculative decoding is a widely adopted technique for accelerating inference in large language models (LLMs), but its performance degrades on long inputs due to increased attention cost and reduced draft accuracy. We introduce SpecExtend, a drop-in enhancement that improves the performance of speculative decoding on long sequences without any additional training. SpecExtend integrates efficient attention mechanisms such as FlashAttention and Hybrid Tree Attention into both the draft and target models, reducing latency across all stages. To improve draft accuracy and speed, we propose Cross-model Retrieval, a novel KV cache update strategy that uses the target model’s attention scores to dynamically select relevant context for the draft model. Extensive evaluations on three long-context understanding datasets show that SpecExtend accelerates standard tree-based speculative decoding by up to 2.22 for inputs up to 16K tokens, providing an effective solution for speculative decoding of long sequences. The code is available at https://github.com/jycha98/SpecExtend.
SpecExtend: A Drop-in Enhancement for Speculative Decoding of
Long Sequences
@@ 1 Introduction
However, the performance of speculative decoding frameworks drops significantly as input length increases. We identify two primary causes: (1) increased latency in both drafting and verification steps due to the quadratic complexity of standard attention, and (2) reduced draft accuracy, as the draft model is typically smaller and trained only on short sequences. Meanwhile, retraining draft models on long contexts is costly, highlighting the need for a drop-in solution that improves long-input performance while preserving the original benefits of existing frameworks.
To this end, we propose SpecExtend, a drop-in enhancement for speculative decoding on long inputs. SpecExtend accelerates the forward passes of both the draft and target models by integrating efficient attention mechanisms across all stages. In order to improve drafting accuracy on long inputs without additional training, we introduce Cross-model Retrieval, a novel cache update strategy for speculative decoding. We dynamically update the draft model’s KV cache with globally relevant context, guided by the target model’s attention scores.
@@ 2.1 Efficient Attention Mechanisms
Unlike prefill, the decoding phase uses cached key and value (KV) states and computes attention only with the newly generated tokens as query. FlashDecoding Dao (2024) accelerates this step by parallelizing across the KV dimension, improving efficiency for short queries. However, it is incompatible with the tree masks required by tree-based speculative decoding frameworks Miao et al. (2024). To resolve this, LongSpec Yang et al. (2025) introduces Hybrid Tree Attention, which splits the KV cache into two parts: the cached segment that requires no masking and the speculative segment that applies the tree mask. FlashDecoding is applied to the cached segment, while standard attention is applied to the speculative part. The outputs are then fused using a log-sum-exp operation, allowing efficient tree attention computation for long inputs.
@@ 2.2 Cross-model Retrieval Cache
We assess the effectiveness of cross-model retrieval using the Needle Retrieval task Li et al. (2024a); Contributors (2023). Specifically, we measure how well the draft model utilizes the retrieved context to locate and draft the tokens of a needle in long inputs. We compare its accuracy against three draft model cache types: (1) Full KV Cache, (2) StreamingLLM Xiao et al. (2023), which retains only the initial and most recent tokens using a static cache policy, and (2) TriForce Sun et al. (2024), which also uses the target model’s attention scores to retrieve top chunks, but employs the same large model for both drafting and verification. While accurate, drafting with the target model is slow due to its large model weights. TriForce serves as an upper bound on how well the retrieved context can be utilized by a smaller draft model.
@@ 3.1 Experiment Setting
We evaluate SpecExtend’s performance on standard speculative decoding baselines using Vicuna-7B Chiang et al. (2023) and LongChat-7B Li et al. (2023) as target models. Draft models include both EAGLE Li et al. (2024c) and off-the-shelf LLMs, Vicuna-68M Yang et al. (2024) and LLaMA-68M Miao et al. (2023). We adopt tree-based drafting with dynamic tree expansion Wang et al. (2025) and use long summarization task, aiming to generate 256 tokens on GovReport Huang et al. (2021), PG-19 Rae et al. (2019), and BookSum Kryściński et al. (2021). All experiments are conducted with a temperature of 0 on 2 A100 80GB GPUs. Further details are provided in Appendix D.
@@ 3.2 Main Results
Figure 2 shows that SpecExtend effectively reduces inference time across all stages of speculative decoding. Meanwhile, Figure 3 shows that the cross-model retrieval cache significantly improves draft accuracy on long inputs, outperforming the static cache policy of StreamingLLM. These improvements lead to consistent speedup gains across all three datasets with both off-the-shelf LLMs and EAGLE as draft models, as shown in Table 2. We provide ablation studies on each component of SpecExtend in Appendix E.2.
@@ 4 Conclusion
While SpecExtend significantly improves the speedup of speculative decoding frameworks, token generation speed still degrades as input length increases. This is primarily due to the inherent growth in attention computation, even when using efficient mechanisms. In particular, the target model’s prefill and decoding steps for verification remain a bottleneck for long inputs, as SpecExtend still operates on the full KV cache for the target model’s forward passes. Nevertheless, SpecExtend effectively extends the range over which speculative decoding frameworks maintain high performance. Additionally, our method does not accelerate existing frameworks to the point of outperforming those specifically trained for long inputs, such as LongSpec. Nevertheless, our proposed cross-model retrieval cache can be integrated into other solutions to provide further speedup, and SpecExtend provides substantial off-the-shelf acceleration when applied to tree-based frameworks.

### v2

@@ Abstract
Speculative decoding is a widely adopted technique for accelerating inference in large language models (LLMs), but its performance degrades on long inputs due to increased attention cost and reduced draft accuracy. We introduce SpecExtend, a drop-in enhancement that improves the performance of speculative decoding on long sequences without any additional training. First, SpecExtend integrates efficient attention mechanisms such as FlashAttention and Hybrid Tree Attention into both the draft and target models. To improve draft accuracy and speed on long inputs without retraining, we propose Cross-model Retrieval, a novel KV cache eviction strategy that uses the target model’s attention scores to dynamically select relevant context for the draft model. Extensive evaluations on three long-context understanding datasets show that SpecExtend accelerates standard tree-based speculative decoding by up to 2.22× for inputs up to 16K tokens, providing an effective solution for speculative decoding of long sequences. Our code is available at github.com/jycha98/SpecExtend.
SpecExtend: A Drop-in Enhancement for
Speculative Decoding of Long Sequences
@@ 1 Introduction
However, the performance of speculative decoding frameworks drops significantly as input length increases, even at moderate lengths (around 2K tokens for Vicuna-7B and EAGLE). We identify two primary causes: (1) increased latency in both drafting and verification steps due to the quadratic complexity of standard attention, and (2) reduced draft accuracy, as the draft model is typically smaller and trained only on short sequences. Meanwhile, naive key-value (KV) cache eviction methods such as StreamingLLM fail to preserve draft accuracy in tasks like Needle in a Haystack or long-form generation, where the draft model must reference earlier context with finer granularity. Furthermore, retraining draft models on long contexts is costly, highlighting the need for a drop-in solution that improves long-input performance while preserving the original benefits of existing frameworks.
To this end, we propose SpecExtend, a drop-in enhancement for speculative decoding on long inputs. SpecExtend accelerates the forward passes of both the draft and target models by integrating efficient attention mechanisms across all stages. To improve drafting speed and accuracy on long inputs without additional training, we introduce Cross-model Retrieval, a novel cache update strategy for speculative decoding. We dynamically update the draft model’s KV cache with globally relevant context, guided by the target model’s attention scores.
@@ 2.1 Efficient Attention Mechanisms
Unlike prefill, the decoding phase uses cached KV states and computes attention only with the newly generated tokens as query. FlashDecoding Dao (2024) accelerates this step by parallelizing across the KV sequence length, improving efficiency for short queries. However, it is incompatible with the tree masks required by tree-based speculative decoding frameworks Miao et al. (2024). To resolve this, LongSpec Yang et al. (2025) introduces Hybrid Tree Attention, which splits the KV cache into two parts: the cached segment that requires no masking and the speculative segment that applies the tree mask. FlashDecoding is applied to the cached segment, while standard attention is applied to the speculative part. The outputs are then fused using a log-sum-exp operation, allowing efficient tree attention computation for long inputs.
@@ 2.2 Cross-model Retrieval Cache
We assess the effectiveness of Cross-model Retrieval using the Needle Retrieval task Li et al. (2024a). Specifically, we measure how well the draft model utilizes the retrieved context to locate and draft tokens of a needle in long inputs. We compare its accuracy against three draft model cache types: (1) Full KV Cache, (2) StreamingLLM Xiao et al. (2023), which retains only the initial and most recent tokens using a static cache policy, and (3) TriForce Sun et al. (2024), which also uses the target model’s attention scores to retrieve top chunks, but employs the same large model for both drafting and verification. While accurate, drafting with the target model is slow due to its large model weights. TriForce serves as an upper bound on how well the retrieved context can be utilized by a smaller draft model.
@@ 3.1 Experiment Setting
We evaluate SpecExtend’s performance on standard speculative decoding baselines using Vicuna-7B Chiang et al. (2023) and LongChat-7B Li et al. (2023) as target models. Draft models include both EAGLE Li et al. (2024c) and off-the-shelf LLMs, Vicuna-68M Yang et al. (2024) and LLaMA-68M Miao et al. (2023). We adopt tree-based drafting with dynamic tree expansion Wang et al. (2025) and use long summarization task, aiming to generate 256 tokens on GovReport Huang et al. (2021), PG-19 Rae et al. (2019), and BookSum Kryściński et al. (2021). All experiments are conducted with a temperature of 0 on 2 A100 80GB GPUs. Further experiment details are provided in Appendix D.
@@ 3.2 Main Results
Figure 2 shows that the Cross-model Retrieval cache substantially improves draft accuracy on long inputs, outperforming the static cache policy of StreamingLLM. Higher draft accuracy reduces the total number of speculative decoding iterations, and combined with efficient attention mechanisms, leads to a significant reduction in inference time across all stages of decoding as shown in Figure 3. As a result, SpecExtend achieves consistent speedup gains across all three datasets with both off-the-shelf LLMs and EAGLE draft models (Table 2). We provide ablation results for each component of SpecExtend in Appendix E.2.
@@ 4 Conclusion
While SpecExtend significantly improves the speedup of speculative decoding frameworks, token generation speed still degrades as input length increases. This is primarily due to the inherent growth in attention computation, even when using efficient mechanisms. In particular, the target model’s prefill and decoding steps remain a bottleneck for long inputs, as speculative decoding operates on the full KV cache for the target model’s forward passes. Nonetheless, SpecExtend effectively extends the range over which speculative decoding frameworks maintain high performance. SpecExtend does not accelerate existing frameworks to the point of surpassing methods trained specifically for long inputs (e.g., LongSpec). However, because it preserves baseline performance, SpecExtend can outperform LongSpec on short inputs when paired with a strong draft model such as EAGLE-3. In addition, it provides substantial off-the-shelf acceleration for long inputs without requiring retraining. Furthermore, our proposed Cross-model Retrieval cache is compatible with other approaches and can be integrated to achieve further speedup.